The volume of an ideal gas (γ = 1.5) is changed adiabatically from 4.00 litres to 3.00 litres. Find the ratio of (a) the final pressure to the initial pressure and (b) the final temperature to the initial temperature.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. γ = 1.5 Since it is an adiabatic process, So PV^γ = const. (a) P base 1V base 1^γ = P base 2V base 2^γ Given V base 1 = 4 L, V base 2 = 3 L, P base 2/P base 1 = ? ⇒ P base 2/P base 1 = (V base 1/V base 2)^γ = (4/3)^1/5 = 1.5396 = 1.54 (b) TV^γ-1 = Const. T base 1/V base 1)^γ-1 = T base 2/V base 2)^γ-1 ⇒ T base 1/T base 2 = (V base 1/V base 2)^γ-1 = (4/3)^0.5 = 1.154